# The consistency strength of determinacy when all sets are universally Baire

### v2

@@ 5.2. Direct limit systems at $\kappa_{i}$
Finally, we argue that there is an embedding $\sigma$ such that $\pi_{E}\upharpoonright\mathcal{M}_{\infty}^{i}=\sigma\circ\pi_{b}^{\mathcal{T}}$. Recall the internal direct limit system giving rise to $\mathcal{M}_{\infty}^{i}$ in $\mathcal{M}$. $\mathcal{M}_{b}^{\mathcal{T}}$ is iterable via a tail of $\Sigma$ in $\mathcal{W}$ and hence strongly $f_{l}^{(i)}$-iterable in $\mathcal{M}$ for every $l<\omega$. Hence $(\mathcal{M}_{b}^{\mathcal{T}},l)\in\mathcal{I}^{(i)}$ for any $l<\omega$. Recall that $\mathcal{M}_{b}^{\mathcal{T}}$ can be obtained as an initial segment of a hybrid backgrounded construction in $\mathcal{M}|\delta_{r}$. By agreement of $\mathcal{M}$ and $\pi_{E}(\mathcal{M})$ up to the strength of $E$ and hence on $\mathcal{M}_{b}^{\mathcal{T}}\unlhd\mathcal{N}$, $\mathcal{M}_{b}^{\mathcal{T}}$ is strongly $\pi_{E}(f_{l}^{(i)})$-iterable in $\pi_{E}(\mathcal{M})$, where $\pi_{E}(f_{l}^{(i)})$ is the canonical function obtained from the images of the terms $\tau_{k}^{(i)}$, $k<l$, under $\pi_{E}$. Let $\sigma_{l}$ be the corresponding direct limit embedding into $\pi_{E}(\mathcal{M}_{\infty}^{i})$. Then $\sigma=\bigcup_{l<\omega}\sigma_{l}\colon\mathcal{M}_{b}^{\mathcal{T}}\rightarrow\pi_{E}(\mathcal{M}_{\infty}^{i})$ is an embedding in $\mathcal{M}$. We are left with showing that $\pi_{E}\upharpoonright\mathcal{M}_{\infty}^{i}=\sigma\circ\pi_{b}^{\mathcal{T}}$.
@@
In the case that $E$ is added by the fully backgrounded part of the $\mathcal{M}$-construction, this follows as in the proof of the successor case in the proof of [12, Theorem 6.5] (see also Lemma 5.16 below). In the case that $E$ is added by the generically countably complete part of the construction, i.e., for the “moreover” part of the claim, the argument for (2) above shows that
$\pi_{E}\upharpoonright\mathcal{M}_{\infty}^{i}=\pi^{\Sigma_{\mathcal{M}_{\infty}^{i}}}_{\mathcal{M}_{\infty}^{i},\pi_{E}(\mathcal{M}_{\infty}^{i})}=\sigma\circ\pi_{b}^{\mathcal{T}}$
by our choice of $\sigma$ and $b=\Sigma_{\mathcal{M}_{\infty}^{i}}(\mathcal{T})$. ∎
@@ 5.3. Certified extenders
Let $F$ be an extender with critical point $\kappa_{i}$ for some $i<\omega$. Then we say $F$ is certified iff
$\pi_{F}\upharpoonright\mathcal{M}_{\infty}^{i,-}=\pi^{\Sigma^{i,\infty}}_{\mathcal{M}_{\infty}^{i,-},\pi_{F}(\mathcal{M}_{\infty}^{i,-})},$
for $\Sigma^{i,\infty}=\Sigma_{\mathcal{M}_{\infty}^{i,-}}$ the tail strategy of $\Sigma_{i}$ as described in the previous subsection. Here $\mathcal{M}_{\infty}^{i,-}$ denotes $\mathcal{M}_{\infty}^{i}|\delta_{\infty}^{i}$, where $\delta_{\infty}^{i}$ denotes the largest Woodin cardinal of the $\kappa_{i}$-suitable premouse $\mathcal{M}_{\infty}^{i}$.
@@
An extender $F$ on $\mathcal{M}$ with critical point $\kappa_{i}$ is uniquely determined by $\pi_{F}\upharpoonright\mathcal{M}_{\infty}^{i,-}$ as $\mathcal{M}_{\infty}^{i,-}\cap\operatorname{Ord}=(\kappa_{i}^{+})^{\mathcal{M}}$ by the following lemma.
@@
Let $i\geq 0$. Then
$\mathcal{M}_{\infty}^{i,-}\cap\operatorname{Ord}=(\kappa_{i}^{+})^{\mathcal{M}}.$
@@
The proof uses similar ideas as [14, Lemma 2.7(b)] (cf., also [29, Lemma 3.38(2)]). For the easier inequality $\mathcal{M}_{\infty}^{i,-}\cap\operatorname{Ord}=\delta_{\infty}^{i}\leq(\kappa_{i}^{+})^{\mathcal{M}}$, let $\alpha<\delta_{\infty}^{i}$. Then there is some $(\mathcal{Q},l)\in\mathcal{I}^{(i)}$ and some $\beta<\gamma_{f_{l}^{(i)}}^{\mathcal{Q}}$ such that $\pi_{(\mathcal{Q},l),\infty}(\beta)=\alpha$. Consider
@@
By $\mathcal{P}$-constructions [^17] See [12, Lemma 3.42] for $\mathcal{P}$-constructions in this hybrid setting. For ordinary mice, $\mathcal{P}$-constructions were introduced by John Steel. This is why Sargsyan calls them $S$-constructions in [12]. in $\mathcal{W}$, (2) implies that there is a proper class $\Sigma$-mouse $\mathcal{P}(\mathcal{N})$ over $\mathcal{N}$ such that $\mathcal{W}|\zeta$ is generic over $\mathcal{P}(\mathcal{N})$ and $\mathcal{P}(\mathcal{N})[\mathcal{W}|\zeta]=\mathcal{W}$.
@@ 5.4. Preservation of certification under hulls
We first argue that taking hulls preserves the certification of extenders in case the critical points of the extenders do not get collapsed.
@@
Let $i<\omega$ and let $\bar{\mathcal{W}}$ be the transitive collapse of a hull of $\mathcal{W}|\Omega$ for some sufficiently large $\Omega$ such that $\mathcal{W}|\kappa_{i}\subseteq\bar{\mathcal{W}}$. Let
@@
be the uncollapse embedding, $\bar{\mathcal{M}}$ the collapse of $\mathcal{M}|\Omega$, and $(\bar{\kappa}_{j},\bar{\delta}_{j}\mid j<\omega)$ the preimages of $(\kappa_{j},\delta_{j}\mid j<\omega)$ (so in particular, $\bar{\kappa}_{j}=\kappa_{j}$ and $\bar{\delta}_{j}=\delta_{j}$ for $j\leq i$). Then every extender in $\bar{\mathcal{M}}$ with critical point $\kappa_{i}$ is certified in $\bar{\mathcal{M}}$.
Remark.
Note that Lemma 5.14 is not trivial even though we do not collapse the direct limit $\mathcal{M}^{i}_{\infty}$. The extender we want to certify gets collapsed as its index gets collapsed, so the certification in the collapse $\bar{\mathcal{W}}$ uses a different tree on $\mathcal{M}^{i}_{\infty}$ than the certification in $\mathcal{W}$ does.
Proof of Lemma 5.14.
Let $E$ be an $\bar{\mathcal{M}}$-extender with critical point $\kappa_{i}$ and let $F=\pi(E)$ be its preimage under the collapse. We aim to show that
$\pi_{E}\upharpoonright\mathcal{M}_{\infty}^{i,-}=\pi^{\Sigma^{i,\infty}}_{\mathcal{M}_{\infty}^{i,-},\pi_{E}(\mathcal{M}_{\infty}^{i,-})},$
for $\Sigma^{i,\infty}=(\Sigma_{i})_{\mathcal{M}_{\infty}^{i,-}}$. As $F$ is certified in $\mathcal{M}$, we can find an iteration tree $\mathcal{U}$ on $\mathcal{M}_{\infty}^{i,-}$ of limit length and a cofinal well-founded branch $b$ through $\mathcal{U}$ such that $\pi_{F}(\mathcal{M}_{\infty}^{i,-})=\mathcal{M}_{b}^{\mathcal{U}}$ and
$\pi^{\Sigma^{i,\infty}}_{\mathcal{M}_{\infty}^{i,-},\pi_{F}(\mathcal{M}_{\infty}^{i,-})}=\pi_{b}^{\mathcal{U}}.$
In particular, $\mathcal{U}^{\frown}b$ is by $\Sigma^{i,\infty}$. Let $\bar{\mathcal{U}}$ and $\bar{b}$ be the images of $\mathcal{U}$ and $b$ under the collapse, i.e., $\pi(\bar{\mathcal{U}},\bar{b})=(\mathcal{U},b)$. Then by hull condensation (cf. Definition 3.1), $\bar{\mathcal{U}}^{\frown}\bar{b}$ is by $\Sigma^{i,\infty}$ as well.
By elementarity of $\pi$, we have $\pi_{E}(\mathcal{M}_{\infty}^{i,-})=\mathcal{M}_{\bar{b}}^{\bar{\mathcal{U}}}$ as $\mathcal{M}_{\infty}^{i,-}$ is fixed by the collapse. So
$\pi_{\bar{b}}^{\bar{\mathcal{U}}}\colon\mathcal{M}_{\infty}^{i,-}\rightarrow\pi_{E}(\mathcal{M}_{\infty}^{i,-}).$
As $\bar{\mathcal{U}}^{\frown}\bar{b}$ is by $\Sigma^{i,\infty}$ this implies
$\pi^{\Sigma^{i,\infty}}_{\mathcal{M}_{\infty}^{i,-},\pi_{E}(\mathcal{M}_{\infty}^{i,-})}=\pi_{\bar{b}}^{\bar{\mathcal{U}}}.$
Note that the fact that $F$ is certified allows us to define $b=\Sigma^{i,\infty}(\mathcal{U})$ via $F$ as follows: $b$ is the unique branch such that
$\pi_{F}\,"\mathcal{M}_{\infty}^{i,-}=\operatorname{rng}(\pi_{b}^{\mathcal{U}}).$
@@
$\pi_{E}\,"\mathcal{M}_{\infty}^{i,-}=\operatorname{rng}(\pi_{\bar{b}}^{\bar{\mathcal{U}}}).$
@@
$\pi_{E}\upharpoonright\mathcal{M}_{\infty}^{i,-}=\pi_{\bar{b}}^{\bar{\mathcal{U}}}.$
@@
Suppose toward a contradiction that $\pi_{E}\upharpoonright\mathcal{M}_{\infty}^{i,-}\neq\pi_{\bar{b}}^{\bar{\mathcal{U}}}$ and let $a$ be minimal (in the canonical well-order of $\mathcal{M}_{\infty}^{i,-}$) with $\pi_{E}(a)\neq\pi_{\bar{b}}^{\bar{\mathcal{U}}}(a)$. Say $\pi_{E}(a)>\pi_{\bar{b}}^{\bar{\mathcal{U}}}(a)$. Then $\pi_{\bar{b}}^{\bar{\mathcal{U}}}(a)\not\in\operatorname{rng}(\pi_{E}\upharpoonright\mathcal{M}_{\infty}^{i,-})=\pi_{E}\,"\mathcal{M}_{\infty}^{i,-}$ as there is no element of $\mathcal{M}_{\infty}^{i,-}$ that $\pi_{E}$ can map to $\pi_{\bar{b}}^{\bar{\mathcal{U}}}(a)$ while respecting the well-order of $\mathcal{M}_{\infty}^{i,-}$. This contradicts $\pi_{E}\,"\mathcal{M}_{\infty}^{i,-}=\operatorname{rng}(\pi_{\bar{b}}^{\bar{\mathcal{U}}})$. The argument in the case $\pi_{E}(a)<\pi_{\bar{b}}^{\bar{\mathcal{U}}}(a)$ is similar. ∎
Now we extend Lemma 5.14 to include hulls that collapse the critical point of the extender in question.

### v3

@@ 5.2. Direct limit systems at $\kappa_{i}$
Finally, we argue that there is an embedding $\sigma$ such that $\pi_{E}\upharpoonright\mathcal{P}_{\infty}^{i}=\sigma\circ\pi_{b}^{\mathcal{T}}$. Recall the internal direct limit system giving rise to $\mathcal{P}_{\infty}^{i}$ in $\mathcal{M}^{(k)}|\delta_{k+1}$. $\mathcal{M}_{b}^{\mathcal{T}}$ is iterable via a tail of $\Sigma$ in $\mathcal{W}$ and hence strongly $f_{l}^{(i)}$-iterable in $\mathcal{M}^{(k)}|\delta_{k+1}$ for every $l<\omega$. Hence $(\mathcal{M}_{b}^{\mathcal{T}},l)\in\mathcal{I}^{(i)}$ for any $l<\omega$. Recall that $\mathcal{M}_{b}^{\mathcal{T}}$ can be obtained as an initial segment of a hybrid backgrounded construction in $\mathcal{M}^{(r)}|\delta_{r}$. By agreement of $\mathcal{M}^{(r)}|\delta_{r+1}$ and $\pi_{E}(\mathcal{M}^{(r)}|\delta_{r+1})$ up to the strength of $E$ and hence on $\mathcal{M}_{b}^{\mathcal{T}}\unlhd\mathcal{N}$, $\mathcal{M}_{b}^{\mathcal{T}}$ is strongly $\pi_{E}(f_{l}^{(i)})$-iterable in $\pi_{E}(\mathcal{M}^{(r)}|\delta_{r+1})$, where $\pi_{E}(f_{l}^{(i)})$ is the canonical function obtained from the images of the terms $\tau_{n}^{(i)}$, $n<l$, under $\pi_{E}$. Let $\sigma_{l}$ be the corresponding direct limit embedding into $\pi_{E}(\mathcal{P}_{\infty}^{i})$. Then $\sigma=\bigcup_{l<\omega}\sigma_{l}\colon\mathcal{M}_{b}^{\mathcal{T}}\rightarrow\pi_{E}(\mathcal{P}_{\infty}^{i})$ is an embedding in $\mathcal{M}^{(r)}|\delta_{r+1}$. We are left with showing that $\pi_{E}\upharpoonright\mathcal{P}_{\infty}^{i}=\sigma\circ\pi_{b}^{\mathcal{T}}$.
@@
In the case that $E$ is added by the fully backgrounded part of the $\mathcal{M}$-construction, this follows as in the proof of the successor case in the proof of [13, Theorem 6.5] (see also Lemma 5.16 below). In the case that $E$ is added by the generically countably complete part of the construction, i.e., in the case $k>i$, the argument for (2) above shows that
$\pi_{E}\upharpoonright\mathcal{P}_{\infty}^{i}=\pi^{\Sigma_{\mathcal{P}_{\infty}^{i}}}_{\mathcal{P}_{\infty}^{i},\pi_{E}(\mathcal{P}_{\infty}^{i})}=\sigma\circ\pi_{b}^{\mathcal{T}}$
by our choice of $\sigma$ and $b=\Sigma_{\mathcal{P}_{\infty}^{i}}(\mathcal{T})$. ∎
@@ 5.3. Certified extenders
Let $F$ be an extender with critical point $\kappa_{i}$ for some $i<\omega$. Then we say $F$ is certified (in $\mathcal{W}$) iff
$\pi_{F}\upharpoonright\mathcal{P}_{\infty}^{i,-}=\pi^{\Sigma^{i,\infty}}_{\mathcal{P}_{\infty}^{i,-},\pi_{F}(\mathcal{P}_{\infty}^{i,-})},$
for $\Sigma^{i,\infty}=\Sigma_{\mathcal{P}_{\infty}^{i,-}}$ the tail strategy of $\Sigma_{i}$ as described in the previous subsection. Here $\mathcal{P}_{\infty}^{i,-}$ denotes $\mathcal{P}_{\infty}^{i}|\delta_{\infty}^{i}$, where $\delta_{\infty}^{i}$ denotes the largest Woodin cardinal of the $\kappa_{i}$-suitable premouse $\mathcal{P}_{\infty}^{i}$.
@@
An extender $F$ on $\mathcal{M}$ with critical point $\kappa_{i}$ is uniquely determined by $\operatorname{Ult}(\mathcal{W},F)|\pi_{F}(\kappa_{i})$ together with $\pi_{F}\upharpoonright\mathcal{P}_{\infty}^{i,-}$ as $\mathcal{P}_{\infty}^{i,-}\cap\operatorname{Ord}=(\kappa_{i}^{+})^{\mathcal{M}^{(i)}}$ by the following lemma. Recall that under $(\mathsf{IH})_{i}$, $\mathcal{M}^{(i)}|\delta_{i+1}=\mathcal{M}|\delta_{i+1}$ is well-defined and well-founded and, in particular, $(\kappa_{i}^{+})^{\mathcal{M}^{(i)}|\delta_{i+1}}=(\kappa_{i}^{+})^{\mathcal{M}^{(i)}}=(\kappa_{i}^{+})^{\mathcal{M}}$. So we will write $(\kappa_{i}^{+})^{\mathcal{M}}$ in this situation to simplify the notation.
@@
Let $i\geq 0$ and suppose $(\mathsf{IH})_{i}$. Then
$\mathcal{P}_{\infty}^{i,-}\cap\operatorname{Ord}=(\kappa_{i}^{+})^{\mathcal{M}}.$
@@
The proof uses similar ideas as [15, Lemma 2.7(b)] (cf., also [31, Lemma 3.38(2)]). For the easier inequality $\mathcal{P}_{\infty}^{i,-}\cap\operatorname{Ord}=\delta_{\infty}^{i}\leq(\kappa_{i}^{+})^{\mathcal{M}}$, let $\alpha<\delta_{\infty}^{i}$. Then there is some $(\mathcal{Q},l)\in\mathcal{I}^{(i)}$ and some $\beta<\gamma_{f_{l}^{(i)}}^{\mathcal{Q}}$ such that $\pi_{(\mathcal{Q},l),\infty}(\beta)=\alpha$. Consider
@@
By $\mathcal{P}$-constructions [^24] See [13, Lemma 3.42] for $\mathcal{P}$-constructions in this hybrid setting. For ordinary mice, $\mathcal{P}$-constructions were introduced by John Steel. This is why Sargsyan calls them $S$-constructions in [13]. in $\mathcal{W}$, (2) implies that there is a proper class $\Sigma$-mouse $\mathcal{P}(\mathcal{N})$ over $\mathcal{N}$ such that $\mathcal{W}|\zeta$ is generic over $\mathcal{P}(\mathcal{N})$ and $\mathcal{P}(\mathcal{N})[\mathcal{W}|\zeta]=\mathcal{W}$.
@@ 5.4. Preservation of certification under hulls
We argue that taking hulls preserves the certification of extenders.
@@
Let $i<\omega$ and suppose $(\mathsf{IH})_{i}$, so $\mathcal{M}^{(i)}|\delta_{i+1}$ is well-defined. Let $\bar{\mathcal{W}}$ be the transitive collapse of a hull of $\mathcal{W}|\Omega$ for some sufficiently large $\Omega$ such that $\mathcal{W}|(\delta_{i}^{+\omega})^{\mathcal{W}}\subseteq\bar{\mathcal{W}}$. Let
@@
be the uncollapse embedding, $\bar{\mathcal{M}}$ the collapse of $\mathcal{M}^{(i)}|\Omega$, and $(\bar{\kappa}_{j},\bar{\delta}_{j}\mid j<\omega)$ the preimages of $(\kappa_{j},\delta_{j}\mid j<\omega)$. [^25] So, in particular, $\bar{\kappa}_{j}=\kappa_{j}$ and $\bar{\delta}_{j}=\delta_{j}$ for $j<i$ and $\bar{\delta}_{i}=\delta_{i}$. Then every extender on the sequence of $\bar{\mathcal{M}}$ with critical point $\kappa_{i}$ is certified in $\bar{\mathcal{W}}$.
Proof.
Let $E$ be an extender on the sequence of $\bar{\mathcal{M}}$ with critical point $\kappa_{i}$ and let $F=\pi(E)$ be its preimage under the collapse. We aim to show that
$\pi_{E}\upharpoonright\bar{\mathcal{P}}_{\infty}^{i,-}=\pi^{\bar{\Sigma}^{i,\infty}}_{\bar{\mathcal{P}}_{\infty}^{i,-},\pi_{E}(\bar{\mathcal{P}}_{\infty}^{i,-})},$
for $\bar{\mathcal{P}}_{\infty}^{i,-}$ the collapse of $\mathcal{P}_{\infty}^{i,-}$ and $\bar{\Sigma}^{i,\infty}=(\Sigma_{i})_{\bar{\mathcal{P}}_{\infty}^{i,-}}$. As $F$ is an extender on the sequence of $\mathcal{M}^{(i)}$ with critical point $\kappa_{i}$, it is certified in $\mathcal{W}$. Therefore, we can find a stack of iteration trees $\vec{\mathcal{U}}$ on $\mathcal{P}_{\infty}^{i,-}$ with last model $\mathcal{M}^{\vec{\mathcal{U}}}_{\infty}=\pi_{F}(\mathcal{P}_{\infty}^{i,-})$ such that
$\pi^{\Sigma^{i,\infty}}_{\mathcal{P}_{\infty}^{i,-},\pi_{F}(\mathcal{P}_{\infty}^{i,-})}=\pi_{\infty}^{\vec{\mathcal{U}}},$
where $\pi_{\infty}^{\vec{\mathcal{U}}}\colon\mathcal{P}_{\infty}^{i,-}\rightarrow\mathcal{M}_{\infty}^{\vec{\mathcal{U}}}$ denotes the iteration embedding along $\vec{\mathcal{U}}$. In particular, $\vec{\mathcal{U}}$ is by $\Sigma^{i,\infty}$. Let $\vec{\mathcal{T}}$ be the stack of iteration trees on $\mathcal{P}^{i,-}$ with last model $\mathcal{P}_{\infty}^{i,-}$ according to $\Sigma^{i}$. Let $\bar{\mathcal{T}}$ and $\bar{\mathcal{U}}$ be the images of $\vec{\mathcal{T}}$ and $\vec{\mathcal{U}}$ under the collapse, i.e., $\pi(\bar{\mathcal{T}},\bar{\mathcal{U}})=(\vec{\mathcal{T}},\vec{\mathcal{U}})$. Then by hull condensation (cf. Definition 3.1) applied to the stacks $\vec{\mathcal{T}}^{\frown}\vec{\mathcal{U}}$ and $\bar{\mathcal{T}}^{\frown}\bar{\mathcal{U}}$ on $\mathcal{P}^{i,-}$, the stack $\bar{\mathcal{T}}^{\frown}\bar{\mathcal{U}}$ is by $\Sigma^{i}$ as well. In particular, $\bar{\mathcal{U}}$ is by $\bar{\Sigma}^{i,\infty}$. By elementarity of $\pi$, we have $\pi_{E}(\bar{\mathcal{P}}_{\infty}^{i,-})=\mathcal{M}_{\infty}^{\bar{\mathcal{U}}}$. So
$\pi_{\infty}^{\bar{\mathcal{U}}}\colon\bar{\mathcal{P}}_{\infty}^{i,-}\rightarrow\pi_{E}(\bar{\mathcal{P}}_{\infty}^{i,-}).$
As $\bar{\mathcal{U}}$ is by $\bar{\Sigma}^{i,\infty}$ this implies
$\pi^{\bar{\Sigma}^{i,\infty}}_{\bar{\mathcal{P}}_{\infty}^{i,-},\pi_{E}(\bar{\mathcal{P}}_{\infty}^{i,-})}=\pi_{\infty}^{\bar{\mathcal{U}}}.$
Note that as $F$ is certified,
$\pi_{F}\,"\mathcal{P}_{\infty}^{i,-}=\operatorname{rng}(\pi_{\infty}^{\vec{\mathcal{U}}}).$
@@
$\pi_{E}\,"\bar{\mathcal{P}}_{\infty}^{i,-}=\operatorname{rng}(\pi_{\infty}^{\bar{\mathcal{U}}}).$
@@
$\pi_{E}\upharpoonright\bar{\mathcal{P}}_{\infty}^{i,-}=\pi_{\infty}^{\bar{\mathcal{U}}}.$
@@
Suppose toward a contradiction that $\pi_{E}\upharpoonright\bar{\mathcal{P}}_{\infty}^{i,-}\neq\pi_{\infty}^{\bar{\mathcal{U}}}$ and let $a$ be minimal (in the canonical well-order of $\bar{\mathcal{P}}_{\infty}^{i,-}$) with $\pi_{E}(a)\neq\pi_{\infty}^{\bar{\mathcal{U}}}(a)$. Say $\pi_{E}(a)>\pi_{\infty}^{\bar{\mathcal{U}}}(a)$. Then $\pi_{\infty}^{\bar{\mathcal{U}}}(a)\not\in\operatorname{rng}(\pi_{E}\upharpoonright\bar{\mathcal{P}}_{\infty}^{i,-})=\pi_{E}\,"\bar{\mathcal{P}}_{\infty}^{i,-}$ as there is no element of $\bar{\mathcal{P}}_{\infty}^{i,-}$ that $\pi_{E}$ can map to $\pi_{\infty}^{\bar{\mathcal{U}}}(a)$ while respecting the well-order of $\bar{\mathcal{P}}_{\infty}^{i,-}$. This contradicts $\pi_{E}\,"\bar{\mathcal{P}}_{\infty}^{i,-}=\operatorname{rng}(\pi_{\infty}^{\bar{\mathcal{U}}})$. The argument in the case $\pi_{E}(a)<\pi_{\infty}^{\bar{\mathcal{U}}}(a)$ is similar. ∎
5.5. All extenders on the $\mathcal{M}$-sequence are certified
Similarly as in [14, Lemma 2.25] we have that fully backgrounded extenders with critical point $\kappa_{i}$ for some $i\geq 0$ are certified.